maxterms minterms 0 1 0 1 0 1 10 101

1- Write the unsimplified POS Boolean equation for F from the Truth Table.

F =

2- Write the unsimplified SOP Boolean equation for F' from the Truth Table.

F' =

3- Using only DeMorgan’s Theorem (show steps) and the unsimplified POS Boolean equation,

find.
Answer #1

pole Given Thange P8 ser change oooooo O B c ol El Maxterms! Il minemy 0 0 0 0 0 0 0 0 1 0 0 0 100 - - - - Oo---0-0-00-0 do.Subject:............. xati.. ..... ») Unsimplified Sop Boolean equation is E = smco, 1, 2, 6, 8, 10, 11, 13) - A B C D + ABCD(A+B+C) f = (A+B+C) (Ate+D) A + B +D) (A + B to că + B+CD) ((A+B+D) + c)(64*13) = A + B + c) (A+E+D) (A + B + c) (A + B + cto